Consider a student loan of ​$17,500 at a fixed APR of 12​% for 25 years.

a. Calculate the monthly payment.

b. Determine the total amount paid over the term of the loan. c.

Of the total amount​ paid, what percentage is paid toward the principal and what percentage is paid for interest.

(Do not round until the final answer. Then round to the nearest cent as​ needed.)
Answer #1

a.

Calculate the monthly payment

=17500/((1-(1+(12%/12))^(-25*12))/(12%/12))

=184.3142 or 184.31

b. Determine the total amount paid over the term of the loan

=(17500/((1-(1+(12%/12))^(-25*12))/(12%/12)))*(25*12)

=55294.27

c.

paid for interest is=55294.27-17500=37794.27

what percentage is paid toward the principal=17500/55294.27=31.65%

what percentage is paid for interest=37794.27/55294.27=68.35%

the above is answer..